OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y − x2 = −8x − 9?

OpenStudy (anonymous):

y-x^2=-8x-9 y=x^2-8x-9 x=-1 and 9

OpenStudy (anonymous):

Put y = 0 and then find x..

OpenStudy (anonymous):

And that will be your x intercept..

OpenStudy (anonymous):

to find the x intercept you must set y=0 so in this case we got x^2 -8x -9=0 and now we have to use the quadratic formula or the product sum method to solve for the xvalues so the product is -9 and sum is -8 our 2 numbers are -9 and 1 so (x-9)(x+1) = 0 so therefore x=9 and x=-1

OpenStudy (anonymous):

\[\huge X-Intercept : Put \quad y = 0\] \[\huge Y-Intercept : Put \quad x = 0\]

OpenStudy (anonymous):

thanks you guys

OpenStudy (anonymous):

Welcome dear..